# XDiag: Exact Diagonalization for Quantum Many-Body Systems

### v2

@@ Abstract
Exact diagonalization (ED) is a cornerstone technique in quantum many-body physics, enabling precise solutions to the Schrödinger equation for interacting quantum systems [1, 2, 3, 4]. Despite its utility in studying ground states, excited states, and dynamical behaviors, the exponential growth of the Hilbert space with system size presents significant computational challenges. We introduce XDiag, an open-source software package designed to combine advanced and efficient algorithms for ED with and without symmetry-adapted bases with user-friendly interfaces. Implemented in C++ for computational efficiency and wrapped in Julia for ease of use, XDiag provides a comprehensive toolkit for ED calculations. Key features of XDiag include the first publicly accessible implementation of sublattice coding algorithms for large-scale spin system diagonalizations, efficient Lin table algorithms for symmetry lookups, and random-hashing techniques for distributed memory parallelization. The library supports various Hilbert space types (e.g., spin-1/2, electron, and t-J models), facilitates symmetry-adapted block calculations, and automates symmetry considerations. The package is complemented by extensive documentation, a user guide, reproducible benchmarks demonstrating near-linear scaling on thousands of CPU cores, and over 20 examples covering ground-state calculations, spectral functions, time evolution, and thermal states. By integrating high-performance computing with accessible scripting capabilities, XDiag allows researchers to perform state-of-the-art ED simulations and explore quantum many-body phenomena with unprecedented flexibility and efficiency.
@@ 1 Introduction
A particular strength of ED is to perform calculations in a symmetry-adapted basis of the Hilbert space. Not only does employing symmetries reduce the resource requirements and speed up calculations, but it also allows extracting physical insights directly by investigating the spectral structure, also known as tower-of-states analysis [25]. Efficient implementations of symmetry-adapted bases in ED, however, require specialized algorithms to be performed optimally. The so-called Lin tables [5] have been proposed early as an efficient algorithm for indexing Hilbert spaces of spin systems with spin conservation and translational symmetry. Recently, these indexing ideas have been significantly improved to arbitrary spin-$S$ by the library DanceQ [20, 26], which also offers highly efficient distributed memory parallelization. Another route for (memory-)efficient use of space group symmetries is the so-called sublattice coding techniques [7, 27, 28], which have been shown to allow for diagonalizations of spin $S=1/2$ of up to $N=50$ particles [28, 29]. Moreover, several works have also successfully constructed full lattice and spin rotational symmetric bases in ED simulations [30, 31]
@@ 3.1 Hilbert spaces
Figure 1: Enumeration of computational basis states of different Hilbert space blocks. (a) Spinhalf block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins. (b) tJ block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins and $n_{\downarrow}=1$downarrow-spins. (c) Electron block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins and $n_{\downarrow}=1$downarrow-spins. The index of a given basis configuration (represented as an object of type ProductState) can be obtained using the index function.
@@ 3.4 States
Quantum states $|\psi\rangle$ are represented in XDiag by using a State object. A state with zero coefficients is created either implicitly by calling the constructor of State with a given block, or explicitly by calling the zero_state function.
@@
The $2$-norm $\parallel|\psi\rangle\parallel_{2}$ and dot product $\langle\psi_{1}|\psi_{2}\rangle$ of states can easily be computed using the norm and dot/dotC functions.
@@ 3.5.2 Time evolution
$|\phi(t)\rangle=e^{-iHt}|\psi_{0}\rangle\quad\text{or}\quad|\eta(t)\rangle=e^{-\tau H}|\psi_{0}\rangle.$ (3)
@@ 3.6 Measurements
$\langle O\rangle=\langle\psi|O|\psi\rangle,$ (4)
can be evaluated using the inner function. For example, we compute a static spin correlation $\langle S_{0}^{z}S_{j}^{z}\rangle$ between site 00 (resp. $1$ in Julia) and $j$.
@@
Notice, that in Julia sites start counting from $1$, whereas in C++ sites are counted from 00. Furthermore, if a complex wave function or operator is involved, the function innerC in C++ should be called, which returns a complex number. In Julia only inner is available whose return type is decided at runtime.
@@ 3.7.1 Reading from TOML
The TOML input files assume that counting starts at 00 for both the C++ and Julia version, such that the same input file can be used for both languages. In Julia, however, the site index is increased by one after reading in the OpSum.
@@ 3.8 Symmetries
Symmetries are fundamental properties of any physical system. In quantum mechanics, symmetries of a Hamiltonian lead to a set of conserved quantities, also referred to as quantum numbers. Mathematically, quantum numbers label irreducible representations of symmetry groups. A particularly important set of symmetries are space group symmetries, such as translation symmetries in a solid or point group symmetries in molecules. Abstractly, these symmetries are permutations of sites of the interaction graph. Permutation groups then have irreducible representations, which can denote the momentum for translation groups or angular momentum in point groups. XDiag can employ these symmetries, allowing for more efficient computations as well as physical insights, e.g., via tower-of-states analyses [55].
@@ 3.8.1 Permutations
Mathematically, a permutation $\pi$ of $N$ elements is a bijective mapping,
$\pi:\{1,2,\ldots,N\}\rightarrow\{1,2,\ldots,N\},$ (5)
@@ 3.8.2 Representations
Mathematically, a representation $\rho$ of a group $G$ is a mapping from a group to the group of invertible matrices $\textrm{GL}(V)$ on a vector space $V$,
$\rho:G\mapsto\textrm{GL}(V),$ (8)
@@
$\rho(g\cdot h)=\rho(g)\cdot\rho(h)\quad\textrm{for all}\quad g,h\in G.$ (9)
The dimension of $V$, i.e., the dimension of the matrices, is referred to as the dimension of the representation. The character $\chi_{\rho}$ of a representation $\rho$ denotes the trace of the representation matrices,
$\chi_{\rho}(g)=\textrm{Tr}[\rho(g)]\in\mathbf{C}.$ (10)
@@
An important example is the cyclic group $C_{N}$ of order $N$ consisting e.g. of translations $\{\textrm{I},T,T^{2},\ldots,T^{(N-1)}\}$, which has $N$ irreducible one-dimensional representations $\rho_{k}$ with associated character functions $\chi_{k}$. Those can be labeled by numbers $k=2\pi l/N$ for $l\in\{1,\ldots,N\}$, such that the characters of the $k$-th irreducible representation (irrep) are,
$\chi_{k}(T^{n})=e^{ikn}.$ (11)
@@ 3.9 Symmetry-adapted blocks
Representations $\rho$ of a permutation symmetry group $G$ can then be used to create symmetry-adapted blocks. Given a computational basis state $|\mathbf{\sigma}\rangle=|\sigma_{1}\sigma_{2}\cdots\sigma_{N}\rangle$, the corresponding symmetry-adapted state is given as,
$|\mathbf{\sigma}_{\rho}\rangle\equiv\frac{1}{N_{\rho,\psi}}\sum_{g\in G}\chi_{\rho}(g)^{*}g|\mathbf{\sigma}\rangle.$ (13)
Here, the action of a permutation $g$ on a product state $|\sigma\rangle$ is defined as,
$g|\mathbf{\sigma}\rangle=g|\sigma_{1}\sigma_{2}\cdots\sigma_{N}\rangle=|\sigma_{g(1)}\sigma_{g(2)}\cdots\sigma_{g(N)}\rangle,$ (14)
and $N_{\rho,\psi}$ denotes the normalization constant. For example, consider a spin $S=1/2$ system on a four-site chain with a fourfold translation symmetry group $C_{4}$, the computational basis state $|\mathbf{\sigma}\rangle=|\downarrow\downarrow\uparrow\uparrow\rangle$ and the irreducible representation with $k=\pi$. Then the symmetry-adapted basis state of $|\mathbf{\sigma}\rangle$ is given as,
$|\mathbf{\sigma}_{\pi}\rangle=\frac{1}{2}\left(|\downarrow\downarrow\uparrow\uparrow\rangle-|\downarrow\uparrow\uparrow\downarrow\rangle+|\uparrow\uparrow\downarrow\downarrow\rangle-|\uparrow\downarrow\downarrow\uparrow\rangle\right).$ (15)
We refer to the states $\{g|\mathbf{\sigma}\rangle\;|\;g\in G\}$ as the orbit of $|\mathbf{\sigma}\rangle$. XDiag represents each basis state by an integer value. To represent a symmetry-adapted basis state, XDiag uses the state in the orbit of $|\mathbf{\sigma}\rangle$ with the minimal associated integer value, which is then called the representative of the orbit. For more details on symmetry-adapted wavefunctions in the context of exact diagonalization we refer to Refs. [3, 56].
@@
Figure 2: Enumeration of computational basis states of symmetry-adapted Hilbert space blocks with translational symmetry in the $k=0$ representation. (a) Spinhalf block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins. (b) tJ block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins and $n_{\downarrow}=1$downarrow-spins. (c) Electron block on $N=4$ sites with $n_{\uparrow}=2$uparrow-spins and $n_{\downarrow}=1$downarrow-spins.
@@ 3.10 Reading symmetries from TOML files
The permutation group is defined as an integer matrix, whose rows correspond to individual permutations. Notice, that we start counting sites from 00 in this case. This holds regardless of whether the C++ version or the Julia version is used. The input TOML files always start counting at 00, in Julia, the symmetry number array is increased by one. The representations are given names, k.zero, k.pihalf, k.pi, and have an associated field characters, which is a vector of complex numbers. Complex numbers are represented by a two-element array. Technically, real representations are also allowed to have one single real number instead.
@@ 3.11 Symmetrized operators
where $O_{A}$ denotes a local operator acting on sites $A=\{a_{1},\ldots,a_{l_{A}}\}$ and $\mathcal{L}$ denotes the lattice. The symmetrization of the operator $O$ with respect to a permutation group $\mathcal{G}=\{\pi_{1},\ldots,\pi_{M}\}$ is then defined as
$O^{\mathcal{G}}=\frac{1}{M}\sum_{A\subseteq\mathcal{L}}\sum_{\pi\in\mathcal{G}}O_{\pi(A)},$ (17)
where $\pi(A)=\{\pi(a_{1}),\ldots,\pi(a_{l_{A}})\}$ denotes the permutated set of sites of the local operator $O_{A}$. Additionally, we can also symmetrize with respect to a representation $\rho$,
$O^{\mathcal{G,\rho}}=\frac{1}{M}\sum_{A\subseteq\mathcal{L}}\sum_{\pi\in\mathcal{G}}\chi_{\rho}(\pi)O_{\pi(A)},$ (18)
where $\chi_{\rho}$ denotes the characters of the representation $\rho$. Such symmetrizations of OpSum objects can be performed using the symmetrize function.
@@ 4.1 Ground states
This example demonstrates how ground state expectation values can be computed for the spin $S=1/2$ Heisenberg chain as in Eq. 1 with periodic boundary conditions. The Hamiltonian is invariant under translations and decomposes into $N$ irreps labeled by lattice momentum $2\pi k/N$, as explained in section 3.8.2. The Lanczos algorithm is run on all momentum sub-blocks. Correlation functions are computed by first symmetrizing operators with respect to the symmetry group $C_{N}$.
@@
$\mathcal{H}=J\sum_{n=0}^{N-1}\left(S^{x}_{n}S^{x}_{n+1}+S^{y}_{n}S^{y}_{n+1}+\Delta S^{z}_{n}S^{z}_{n+1}\right).$ (20)
The ground state $|\psi_{0}\rangle$ is obtained using the Lanczos algorithm for a given value of $\Delta$, and functions are provided to construct the reduced density matrix related to the region with the first $\ell$ spins. This is obtained by tracing out the complementary degrees of freedom,
$\rho_{\ell}=\text{Tr}_{\bar{\ell}}\left(|\psi_{0}\rangle\langle\psi_{0}|\right),$ (21)
which in practice is done by looping over all possible spin configurations of the two regions, as described in subsection 3.1. The entanglement entropy is obtained through the eigenvalues of the reduced density matrix, $\lambda_{\alpha}$,
$S_{\ell}=-\text{Tr}\left(\rho_{\ell}\ln\rho_{\ell}\right)=-\sum_{\alpha}\lambda_{\alpha}\ln\lambda_{\alpha}.$ (22)
@@
$\mathcal{H}=-K\sum_{\langle i,j\rangle\parallel\tilde{\gamma}}S_{i}^{\gamma}S_{j}^{\gamma}+J\sum_{\langle i,j\rangle}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j},$ (23)
@@
We compute charge correlations in the ground state of the attractive 2D Hubbard model, at half-filling, to show the $(\pi,\pi)$-charge density wave.
@@ 4.2 Full exact diagonalization
$\mathcal{H}=\frac{1}{\sqrt{N}}\sum_{i\neq j=0}^{N-1}t_{ij}Pc^{\dagger}_{i\alpha}c_{j\alpha}+\frac{1}{\sqrt{N}}\sum_{i<j=0}^{N-1}J_{ij}\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j},$ (24)
where $P$ is the projection on non-doubly occupied sites, $\boldsymbol{S}_{i}=\frac{1}{2}c^{\dagger}_{i\alpha}\boldsymbol{\sigma}_{\alpha\beta}c_{j\beta}$ is the spin operator on site $i$. Both the hoppings $t_{ij}=t^{\ast}_{ji}$ and the exchange interaction $J_{ij}$ are independent random numbers with zero mean and variance $\bar{t^{2}}$ and $\bar{J^{2}}$, respectively. This type of system exhibits a transition from a spin glass to a disordered Fermi liquid phase with increasing doping, as shown in Ref. [38].
@@
$\mathcal{H}=-t\sum_{\langle i,j\rangle,\sigma}Pc^{\dagger}_{i\alpha}c_{j\alpha}+J\sum_{\langle i,j\rangle}\left(\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-\frac{1}{4}n_{i}n_{j}\right)$ (25)
with $J/t=0.4$. We consider an $11$-site triangular lattice (arranged in a 4-4-3 geometry) with one hole in the system. The number of electrons is thus $N-1$ (with equal numbers of spin-up and spin-down electrons).
@@
$H=\sum_{i}J_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}-\sum_{i}(h_{i}\sigma_{i}^{z}+\gamma_{i}\sigma_{i}^{x}),$ (28)
where we take $J_{i}\in[0.8,1.2]$, $h_{i}\in[-W,W]$, $\gamma_{i}=1$, and $W$ is the disorder strength. This model is known to display many-body localization for large enough $W$, i.e, it hosts an extensive set of emergent (local) conserved quantities, often called $l$-bits. To study this, we compute the inverse participation ratio, $\mathrm{IPR}=\sum_{i}|\psi_{i}|^{4}$, most often used to study Anderson localization, for some of the low-lying eigenstates. Additionally, we also study the level spacing ratio, $r=\text{min}(\delta_{n},\delta_{n+1})/\text{max}(\delta_{n},\delta_{n+1})$, where $\delta_{n}=E_{n+1}-E_{n}$.
4.3 Tower of states
Symmetry-breaking $\alpha\text{XX}$ chain
We perform a tower of states (TOS) analysis of the $\alpha\text{XX}$ model [57]. This $S=1/2$ model consists of an array of $N$ sites with long-range XY couplings that decay as a power law with exponent $\alpha$,
$\mathcal{H}=-\dfrac{J}{2}\sum_{i<j}\dfrac{S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j}}{|i-j|^{\alpha}}.$ (29)

### v3

@@ Abstract
Exact diagonalization (ED) is a cornerstone technique in quantum many-body physics, enabling precise solutions to the many-body Schrödinger equation for interacting quantum systems [1, 2, 3, 4]. Despite its utility in studying ground states, excited states, and dynamical behaviors, the exponential growth of the Hilbert space with system size presents significant computational challenges. We introduce XDiag, an open-source software package designed to combine advanced and efficient algorithms for ED with and without symmetry-adapted bases with user-friendly interfaces. Implemented in C++ for computational efficiency and wrapped in Julia for ease of use, XDiag provides a comprehensive toolkit for ED calculations. Key features of XDiag include the first publicly accessible implementation of sublattice coding algorithms for large-scale spin system diagonalizations, efficient Lin table algorithms for symmetry lookups, and random-hashing techniques for distributed memory parallelization. The library supports various Hilbert space types (e.g., spin-1/2, electron, and t-J models), facilitates symmetry-adapted block calculations, and automates symmetry considerations. The package is complemented by extensive documentation, a user guide, reproducible benchmarks demonstrating near-linear scaling on thousands of CPU cores, and over 20 examples covering ground-state calculations, spectral functions, time evolution, and thermal states. By integrating high-performance computing with accessible scripting capabilities, XDiag allows researchers to perform state-of-the-art ED simulations and explore quantum many-body phenomena with unprecedented flexibility and efficiency.
@@ 1 Introduction
A particular strength of ED is to perform calculations in a symmetry-adapted basis of the Hilbert space. Not only does employing symmetries reduce the resource requirements and speed up calculations, but it also allows extracting physical insights directly by investigating the spectral structure, also known as tower-of-states (TOS) analysis [25]. Efficient implementations of symmetry-adapted bases in ED, however, require specialized algorithms to be performed optimally. The so-called Lin tables [5] have been proposed early as an efficient algorithm for indexing Hilbert spaces of spin systems with spin conservation and translational symmetry. Recently, these indexing ideas have been significantly improved to arbitrary spin-$S$ by the library DanceQ [20, 26], which also offers highly efficient distributed memory parallelization. Another route for (memory-)efficient use of space group symmetries is the so-called sublattice coding techniques [7, 27, 28], which have been shown to allow for diagonalizations of spin $S=1/2$ of up to $N=50$ particles [28, 29]. Moreover, several works have also successfully constructed full lattice and spin rotational symmetric bases in ED simulations [30, 31]
@@ 3.1 Hilbert spaces
Figure 1: Enumeration of computational basis states of different Hilbert space blocks. (a) Spinhalf block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins. (b) tJ block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins and $n_{\downarrow}=1$ $\downarrow$-spins. (c) Electron block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins and $n_{\downarrow}=1$ $\downarrow$-spins. The index of a given basis configuration (represented as an object of type ProductState) can be obtained using the index function.
@@ 3.2 Operators
3.3 Matrix type
C++
auto sx = arma::mat({{0, 1},{1, 0}});
auto op = Op("Matrix", 0, sx);
C++
auto sx = arma::mat({{0, 1},{1, 0}});
auto sz = arma::mat({{0.5, 0},{0, -0.5}});
arma::mat sxsz = arma::kron(sx, sz);
arma::mat sxszsxsz = arma::kron(sxsz, sxsz);
auto op_sxsz = Op("Matrix", {0, 1}, sxsz);
auto op_sxszsxsz = Op("Matrix", {0, 1, 2, 3}, sxsz);
Here, we have used the Kronecker product function kron.
3.4 Complex couplings
  1. 1.
  2. 2.
[…]
@@ 3.6 States
[…]
@@
[…]
@@ 3.7.2 Time evolution
[…]
@@ 3.8 Measurements
[…]
can be evaluated using the inner function. For example, we compute a static spin correlation $\langle S_{0}^{z}S_{j}^{z}\rangle$ between site 0 (resp. $1$ in Julia) and $j$.
@@
Notice, that in Julia sites start counting from $1$, whereas in C++ sites are counted from 0. Furthermore, if a complex wave function or operator is involved, the function innerC in C++ should be called, which returns a complex number. In Julia only inner is available whose return type is decided at runtime.
@@ 3.9.1 Reading from TOML
The TOML input files assume that counting starts at 0 for both the C++ and Julia version, such that the same input file can be used for both languages. In Julia, however, the site index is increased by one after reading in the OpSum.
@@ 3.10 Symmetries
Symmetries are fundamental properties of any physical system. In quantum mechanics, symmetries of a Hamiltonian lead to a set of conserved quantities, also referred to as quantum numbers. Mathematically, quantum numbers label irreducible representations of symmetry groups. A particularly important set of symmetries are space group symmetries, such as translation symmetries in a solid or point group symmetries in molecules. Abstractly, these symmetries are permutations of sites of the interaction graph, which can be efficiently made use of in ED simulations, see e.g.  [55]. Permutation groups then have irreducible representations, which can denote the momentum for translation groups or angular momentum in point groups. XDiag can employ these symmetries, allowing for more efficient computations as well as physical insights, e.g., via tower-of-states analyses [56]. Internally, XDiag checks for consistency of symmetries and representations and allows the use of permutational symmetries also for multi-site interactions.
@@ 3.10.1 Permutations
Mathematically, a permutation of $N$ elements is a bijective mapping,
[…]
@@ 3.10.2 Representations
Mathematically, a representation of a group $G$ is a mapping from a group to the group of invertible matrices $\textrm{GL}(V)$ on a vector space $V$,
[…]
@@
[…]
The dimension of $V$, i.e., the dimension of the matrices, is referred to as the dimension of the representation. The character of a representation denotes the trace of the representation matrices,
[…]
@@
[…]
$_{k}(T^{n})=e^{ikn}.$ (11)
@@ 3.11 Symmetry-adapted blocks
[…]
@@
Figure 2: Enumeration of computational basis states of symmetry-adapted Hilbert space blocks with translational symmetry in the $k=0$ representation. (a) Spinhalf block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins. (b) tJ block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins and $n_{\downarrow}=1$ $\downarrow$-spins. (c) Electron block on $N=4$ sites with $n_{\uparrow}=2$ $\uparrow$-spins and $n_{\downarrow}=1$ $\downarrow$-spins.
@@ 3.12 Reading symmetries from TOML files
The permutation group is defined as an integer matrix, whose rows correspond to individual permutations. Notice, that we start counting sites from 0 in this case. This holds regardless of whether the C++ version or the Julia version is used. The input TOML files always start counting at 0, in Julia, the symmetry number array is increased by one. The representations are given names, k.zero, k.pihalf, k.pi, and have an associated field characters, which is a vector of complex numbers. Complex numbers are represented by a two-element array. Technically, real representations are also allowed to have one single real number instead.
@@ 3.13 Symmetrized operators
where $O_{A}$ denotes a local operator acting on sites $A=\{a_{1},\ldots,a_{l_{A}}\}$ and $\mathcal{L}$ denotes the lattice. The symmetrization of the operator $O$ with respect to a permutation group $\mathcal{G}=\{{}_{1},\ldots,{}_{M}\}$ is then defined as
[…]
where denotes the characters of the representation . Such symmetrizations of OpSum objects can be performed using the symmetrize function.
@@
3.14 Sparse Matrix Capabilities
Working with many-body quantum systems often involves matrices which only have a small number of non-zero elements, also known as sparse matrices. Since storing and handling such objects in the conventional way (i.e. element by element) is very inefficient, XDiag includes basic implementations of three common sparse-matrix types: the coordinate (COO), the compressed-sparse-row (CSR), and the compressed-sparse-column (CSC) formats.
Just as the matrix function can be used to obtain the full matrix representing a given operator ops (in the form of an OpSum object) on a given Hilbert space block (also see Sec. 3.5), there are functions coo_matrix, csr_matrix, and csc_matrix to obtain the same matrix in the respective sparse format.
C++
auto coo_mat = coo_matrix(ops, block);
auto csr_mat = csr_matrix(ops, block);
auto csc_mat = csc_matrix(ops, block);
Julia
coo_mat = coo_matrix(ops, block)
csr_mat = csr_matrix(ops, block)
csc_mat = csc_matrix(ops, block)
Note that the C++ implementation distinguishes between real and complex matrices, e.g., there are the csr_matrix and the csr_matrixC functions. The objects returned by these functions are “raw” in the sense that they are not an instance of a sparse-matrix implementation of another library, but contain all the information to call the respective sparse-matrix constructor of your sparse-matrix library of choice. For instance, the output of csc_matrix can be used to construct the sp_mat type implemented by the Armadillo library in C++ or the SparseMatrixCSC type implemented by SparseArrays (part of the standard library) in Julia.
C++
auto colptr = arma::conv_to<arma::uvec>::from(csc_mat.colptr);
auto row = arma::conv_to<arma::uvec>::from(csc_mat.row);
auto A = arma::sp_mat(colptr, row,
csc_mat.data, csc_mat.nrows, csc_mat.ncols);
Julia
using SparseArrays
jmat = SparseMatrixCSC(csc_mat.nrows, csc_mat.ncols,
csc_mat.colptr, csc_mat.row, csc_mat.data)
While the COO, CSR, and CSC formats are supported for extracting sparse matrices from XDiag, only the CSR format is used internally because it is the only one suitable for parallelized matrix-vector multiplications. The user guide of our online documentation [46] compares the default (matrix-free, i.e., on-the-fly) implementation of Lanczos iterations to using the CSR-based version after building and converting the full Hamiltonian into the CSR format. Based on these examples, at medium system sizes the runtime of both approaches is comparable, while the matrix-free variant will always require substantially less memory. Note that the computational effort and memory required to construct the CSR representation grows significantly with system size such that the matrix-free default implementation will generally run faster on larger systems. However, a clear advantage of sparse-matrix implementations can be expected when sufficient memory is available for the CSR matrix to be pre-computed once before it needs to be reused frequently.
@@ 4.1 Ground states
[…]
@@
[…]
The ground state $|{}_{0}\rangle$ is obtained using the Lanczos algorithm for a given value of , and functions are provided to construct the reduced density matrix related to the region with the first $\ell$ spins. This is obtained by tracing out the complementary degrees of freedom,
[…]
which in practice is done by looping over all possible spin configurations of the two regions, as described in subsection 3.1. The entanglement entropy is obtained through the eigenvalues of the reduced density matrix, ,
$S_{\ell}=-\text{Tr}\left({}_{\ell}\ln{}_{\ell}\right)=-\sum\ln.$ (22)
@@
[…]
@@
[…]
@@ 4.2 Full exact diagonalization
[…]
@@
[…]
with $J/t=0.4$. We consider an $11$-site triangular lattice (arranged in a 4-4-3 geometry) with one hole in the system. The number of electrons is thus N1N-1 (with equal numbers of spin-up and spin-down electrons).
@@
[…]
Lindblad Super-operator spectrum
We obtain the spectrum of the Lindblad super-operator governing the time-evolution of an $S=1/2$ XXZ chain with bulk dephasing [58],
[…]
where $L_{i}$ is the quantum jump operator associated with a Markovian bath acting on site $i$ and the strength of the dissipation. The spectrum is obtained by constructing the vectorized Lindblad operator,
[…]
where $\tilde{S}$ are auxiliary spin degrees of freedom.
The Lindblad spectrum reveals the system’s excitations and decay timescales. A key quantity is the dissipative gap, defined as the absolute value of the real part of the eigenvalue closest to zero. This gap sets the slowest relaxation time, and its dependence on the model parameters plays a central role in the study of dissipative phase transitions [59, 60].
4.3 Tower-of-states
Symmetry-breaking XX chain
We perform a tower-of-states analysis of the XX model [61]. This $S=1/2$ model consists of an array of $N$ sites with long-range XY couplings that decay as a power law with exponent ,
$\mathcal{H}=-\dfrac{J}{2}\sum_{i<j}\dfrac{S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j}}{|i-j|}.$ (31)